Rank the following three stocks bu their total risjk level highest to lowest. Night Ryder has an average return of 10 percent and standard deviation of 38 percent. The average return and standard deviation of WholeMart are 11 percent and 31 percent and of Fruit Fly are 16 percent and 45 percent.

Rank Stock
1.
2.
3.
Answer #1

Risk is measured through Standard Deviation in Finance.

Hence, if the requirement is to rank only by the respective risk levels, Stock having the highest Standard Deviation will be ranked number 1 and the one with lowest Standard Deviation will rank bottom.

Hence,

Rank I : Fruit Fly

Rank II: Night Ryder

Rank III: WholeMart

However, if the objective is to measure risk as a % of return, then we need to use Sharpe ratio.

Sharpe Ratio is calculated in the following way:-

Sharpe Ratio = (Expected Return of the Stock - Risk free Rate of Return) / Standard Deviation of the Stock

Since, the problem doesn't give any information pertaining to risk-free rate of return, let us assume it to be 3%

Sharpe Ratio of Night Ryder :-

(10 - 3) / 38 = 0.184

Sharpe Ratio of WholeMart :-

(11 - 3) / 31 = 0.258

Sharpe Ratio of Fruit Fly :-

(16 - 3) / 45 = 0.289

Hence, the Stock having a higher Sharpe Ratio is the preferred Investment Choice.